Help I'm lost! I need to solve using the quadratic formula. An airplane makes a 990km flight with a tailwind and returns, flying into the same wind. The total flying time is 3 hours 20 minutes, and the airplane's speed in still air is 600 km/h. What is the spped of the wind?
Answer by scott8148(6628)

time = distance / rate

working in hrs ___ 10/3 = [990 / (600 + w)] + [990 / (600 - w)]

multiplying by 3(360000-w^2) the LCD ___ 3600000 - 10w^2 = 3[990(600-w)] + 3[990(600+w)]

0 = w^2 - 3600

w = ±60